OpenStudy (finn):

8. Density is defined as mass per unit volume, or D = m/v . Solve the density equation for the variable v: A. v = D + m B. v = m/D C. v = D/m D. v = D • m

OpenStudy (anonymous):

B

OpenStudy (anonymous):

cross multiplication

OpenStudy (anonymous):

well u take the amt of water as x and alcohol as 35-x (x/35-x)=4/1 so x = 28 L